OpenStudy (anonymous):

First derivative of x/x^2+1 I am confused with the algebraic simplification. Should I use quotient rule?

OpenStudy (anonymous):

I'm assuming you mean x/(x^2+1)? Make sure to use brackets! Yes you should use quotient rule. f=x, g=x^2+1

OpenStudy (anonymous):

ok I get 1(x^2+1)-x(2x)/(x^2+1)^2 ???

OpenStudy (anonymous):

Brackets! (1(x^2+1)-x(2x))/(x^2+1)^2

OpenStudy (anonymous):

ok this next part I can't do correctly

OpenStudy (anonymous):

Simplifying it? Just expand the top of the fraction and combine common terms.

OpenStudy (anonymous):

I can't do it Ive tried believe me lol

OpenStudy (anonymous):

In my eyes the 1 distributes into the (x^2+1)?

OpenStudy (anonymous):

1(x^2+1)-x(2x) =x^2+1-2x^2 =-x^2+1

OpenStudy (anonymous):

Yes, the 1 distributes.

OpenStudy (anonymous):

Ok so the whole thing reduces top and bottom to -x^2+1 right?

OpenStudy (anonymous):

No, just the top. The bottom doesn't change.

OpenStudy (anonymous):

x= +/- 1 ?

OpenStudy (anonymous):

Are you trying to solve for x when it is equals 0 or something now?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

one of the x^2+1 cancels?

OpenStudy (anonymous):

So now you have (-x^2+1)/(x^2+1)^2 right? To solve for when it equals 0, you only need to solve for when the NUMERATOR equals 0. The denominator can be ignored. so solve for -x^2+1=0 You should get x=+-1, so yeah you're right.

OpenStudy (anonymous):

oh crap I think that is the first time I heard its numerator only. Thanks alot!

OpenStudy (anonymous):

Think about it this way: \[\frac{ -x^2+1 }{ (x^2+1)^2 }=0\]Multiply both sides by (x^2+1)^2. You get \[-x^2+1=0\]So you can always multiply both sides by the denominator and it just goes away.

OpenStudy (anonymous):

Ahhhh ok that makes sense